OpenStudy (kinggeorge):

Without using any kind of derivatives, find \[\lim_{x\to\infty}x-\log_{10}(x).\]

OpenStudy (anonymous):

What if you do \[ \left(\frac 1{10}\right)^{x−\log_{10}(x)} \]

OpenStudy (anonymous):

Hmmm...

OpenStudy (anonymous):

\[ 10^{\log_{10}(x)-x}=\frac{x}{10^x} \]

OpenStudy (anonymous):

can you use logic?

OpenStudy (kinggeorge):

This question arose in a calc 1 class at a point where the students have not yet learned derivatives. We could probably use some basic logic, but not much more.

OpenStudy (anonymous):

unless i am missing something, we know log grows slower than any positive power of \(x\) much slower

OpenStudy (anonymous):

put \(x=10^{12}\) get \(10^{12}-12\)

OpenStudy (anonymous):

Hmm, I wonder if there is a clever way to factor out \(x\) from \(10^x\)

OpenStudy (kinggeorge):

That's all fine and dandy, but how do we know it grows slow enough? If all I know about logs are the basic properties such as \(\log(ab)=\log(a)+\log(b)\), and that it goes to infinity as x goes to infinity, I don't see how to do it.

OpenStudy (anonymous):

Hmmm, what about changing the base of the logarithm.

OpenStudy (kinggeorge):

I suppose we could make an argument towards the fact that \(\log(x^2)=\log(x)+\log(x)\), and try and get some kind of inequality about \(x^2\) and \(\log(x)+\log(x)\). You're welcome to try that. It ultimately shouldn't matter much.

OpenStudy (anonymous):

Hmm \[ x>N\implies x-\log_{10}(x)>M \]

OpenStudy (anonymous):

We know it goes to infinity, so this is ultimately what we need to show.

OpenStudy (anonymous):

It is unfair to say that a student would suspect it went to infinity, and then tried to prove it?

OpenStudy (kinggeorge):

The closest I've come (in my opinion), is by assuming the fact that \[x\ge2\log_{10}(x)\]for \(x\) sufficiently large. Unfortunately, proving this inequality requires derivatives, unless I'm missing something clever.

OpenStudy (anonymous):

Wouldn't it be for any arbitrary constant, not just 2, to show sufficiently large?

OpenStudy (kinggeorge):

If we're given that \(a\ge2\log_{10}(a)\) for some \(a\), how do we know it's true for every \(b>a\)?

OpenStudy (anonymous):

Okay how about this. \[\begin{split} x-\log_{10}(x)&>M\\ 10^{x-\log_{10}(x)}&>10^M\\ \frac{10^x}{x}&>10^M\\ \frac{2^x}{x}\times 5^x&>10^M\\ \end{split} \]We know that \(2^x>x\) so \(2^x/x>1\)\[ \frac{2^x}{x}\times 5^x> 5^x>10^M \]

OpenStudy (anonymous):

Also \[ 10^M=5^{M\log_5{10}} \]

OpenStudy (anonymous):

\[ x>N\implies 5^x>5^{M\log_5(10)} \]

OpenStudy (kinggeorge):

My only question about that, is can you show that \(2^x>x\) for \(x\) sufficiently large?

OpenStudy (anonymous):

> is a lot easier to deal with than >>

OpenStudy (anonymous):

Let me think about it...

OpenStudy (anonymous):

Okay, so consider \(g(x)=2^x-x\).\[ \forall x_1,x_2>100\\ x_2>x_1\\ 2^{x_2}>2^{x_1}\\ 2^{x_2}-x_2>2^{x_1}-x_1 \]This means that \(g(x)\) is always increasing.

OpenStudy (anonymous):

It shows:\[ x_2>x_1\implies g(x_2)>g(x_1) \]Meaning \(g(x)\) is an increasing function for arbitrarily large \(x>100\)

OpenStudy (anonymous):

Not only is it always increasing, but it is positive beyond \(100\).

OpenStudy (anonymous):

So if \(2^x\) is greater than \(x\) at 100 and beyond 100 the gap never narrows, then that means for arbitrarily large \(x\) we know \(2^x>x\)

OpenStudy (anonymous):

Is it still a bit iffy?

OpenStudy (kinggeorge):

It's looking better. Give me a moment to verify everything.

OpenStudy (anonymous):

A lot of the inequality stuff only holds for arbitrarily large \(x\). Not for small or negative \(x\) obviously.

OpenStudy (kinggeorge):

Since we're looking at \(x\to\infty\), we can disregard small or negative \(x\). Can you explain how you concluded that the gap never narrows?

OpenStudy (kinggeorge):

Actually, I don't think that matters. We just need \(2^x>x\) for sufficiently large \(x\), and I think this does it.

OpenStudy (anonymous):

The gap between \(2^x\) and \(x\) is \(2^x-x\).

OpenStudy (anonymous):

This gap has to be \(0\) for one to overtake the other.

OpenStudy (anonymous):

Otherwise whoever is greater will retain supremacy.

OpenStudy (kinggeorge):

Excellent work. I think this proves the limit without resorting to derivatives.